[figure: 9.7 cm]

METASTABLE MAGNETIC STRUCTURES IN THE BLUME-CAPEL MODEL

K. Kułakowski^{a}, M. Antoniuk^{a} and A. del Moral^{b} aFaculty of Physics and Nuclear Techniques University of Mining and Metallurgy al. Mickiewicza 30, 30059 Cracow, Poland bLaboratorio de Magnetismo de Solidos Departamento de Fisica de Materia Condensada Universidad de Zaragoza, 50009 Zaragoza, Spain

Recently [1] we have formulated a criterion of stability of magnetic phase formed out during the first-order phase transition from non-magnetic phase at T=0. This criterion was applied [2] to the antiferromagnetic structure of the C15 atomic lattice. In the present paper results of new calculations for a spontaneously formed magnetic phase and its magnetic energy are given. Dynamic phase diagram is constructed for the phase transition. The results are discussed in the context of the pressure-driven phase transition observed experimentally in YMn_{2} [3].

[1] M. Antoniuk and K. Kułakowski, Physica A 242 (1997) 219.

[2] K. Kułakowski, M. Antoniuk, J. Deskur and P. Gawronski, J. Magn. Magn. Mat. (1999), in print.

[3] K. Kułakowski and A. del Moral, Phys. Rev. B 56 (1997) 8928.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : K.Kulakowski

Address for correspondence : Faculty of Physics and Nuclear Techniques University of Mining and Metallurgy al.Mickiewicza 30, 30-059 Cracow, Poland

Email address : kulakowski@novell.ftj.agh.edu.pl